### v1

# On estimation of weighted cumulative residual Tsallis entropy

Abstract
Recently, weighted cumulative residual Tsallis entropy has been introduced in the literature as a generalization of weighted cumulative residual entropy. We study some new properties of weighted cumulative residual Tsallis entropy measure. Next, we propose some non-parametric estimators of this measure. Asymptotic properties of these estimators are discussed. Performance of these estimators are compared by mean squared error. Non-parametric estimators for weighted cumulative residual entropy measure are also discussed. Two uniformity tests are proposed based on an estimator of these two measures and power of the tests are compared with some popular tests. The tests perform reasonably well.
Keywords–asymptotic normality, consistency, order statistics, uniformity test
Mathematics Subject Classification–62B10, 62G05, 94A17
1 Introduction
Entropy is by far regarded as the most important concept in information theory which was first mathematically formulated by Shannon (1948) independently of Wiener (1948). For a discrete random variable (rv) $X$ that takes values $x_{1},x_{2},\cdots,x_{n}$} with probabilities $\mathbf{P}$ = $p_{1},p_{2},\cdots,p_{n}$}, Shannon entropy is defined as
$H(\mathbf{P})=-\sum_{i=1}^{n}p_{i}\log p_{i}.$ (1)
For a continuous rv $X$, it is defined as
$H(X)=-\int_{A}f(x)\log f(x)dx=E\left[-\log f(X)\right],$ (2)
where log is the natural logarithm, $0\log 0=0$ for computational convenience, $f$ is the probability density function (pdf) of $X$, AA is the support of $X$ and $E$ is the expectation operator. It is also known as the differential entropy. It is considered as a measure of uncertainty and represents the average information contained in the underlying rv. Shannon’s work on entropy gave rise to a new branch of applied probability with useful applications in a variety of fields such as economics, thermodynamics, statistical mechanics, mathematical biology, signal processing, statistics and reliability. An enormous amount of research has been carried out over the years by many scholars on generalizations of Shannon entropy and their applications. One of the important generalizations is due to Tsallis (1988), which is known as generalized Tsallis entropy and is given by
$T_{\alpha}(X)=\frac{1}{\alpha-1}\left(1-\int_{0}^{+\infty}f^{\alpha}(x)dx\right),\;0<\alpha\neq 1.$ (3)
The parameter $\alpha$ is called the generalization parameter and when $\alpha\to 1$, $T_{\alpha}(X)\to H(X)$.
Entropy defined in Eq. (1) is always positive but differential entropy in Eq. (2) may be negative. Differential entropy has some drawbacks such as it cannot be defined for distributions that do not have densities. For discrete rvs $X$ and $Y$, conditional entropy of $X$ given $Y$ is zero if and only if (iff) $X$ is a function of $Y$. However, differential conditional entropy of $X$ given $Y$ is zero does not necessarily imply that $X$ is a function of $Y$. Rao et al. (2004) extended differential entropy by replacing the densities with the survival function of the rv. This measure is called cumulative residual entropy (CRE) and for a non-negative continuous rv $X$, it is defined as
$\xi(X)=-\int_{0}^{+\infty}\bar{F}(x)\log\bar{F}(x)dx,$ (4)
where $\bar{F}$ is the survival function of $X$. The CRE overcomes the above mentioned challenges that differential entropy faces and it possesses fundamental properties that Shannon entropy has. Like Shannon entropy, CRE is always non-negative and it increases by adding independent components (observations) and decreases by conditioning. Apart from that, CRE also possesses some useful properties such as it can be defined for both discrete and continuous rvs and can be easily estimated from data using empirical distribution function. Also CRE(X|Y)=0(X|Y)=0 iff $X$ is a function of $Y$ where CRE(X|Y)(X|Y) is the conditional CRE of $X$ given $Y$.
Ever since its introduction, various generalizations of CRE, analogous to the generalizations of Shannon entropy, have been considered by many authors. Sati and Gupta (2015) proposed one such generalization of the Tsallis entropy which is known as cumulative residual Tsallis entropy (CRTE) of order $\alpha$ and is defined as
$T_{\alpha}(X)=\frac{1}{\alpha-1}\left(1-\int_{0}^{+\infty}\bar{F}^{\alpha}(x)\right)dx,\;0<\alpha\neq 1.$ (5)
Rajesh and Sunoj (2019) proposed an alternative version of CRTE as
$\xi_{\alpha}(X)=\frac{1}{\alpha-1}\int_{0}^{+\infty}\left(\bar{F}(x)-\bar{F}^{\alpha}(x)\right)dx,\;0<\alpha\neq 1.$ (6)
Both $T_{\alpha}(X)$ and $\xi_{\alpha}(X)$ reduce to CRE as $\alpha\to 1$. When $\alpha=2$, $\xi_{\alpha}(X)$ becomes Gini’s mean difference (GMD). This alternative measure is more flexible than $T_{\alpha}(X)$ because it has more relationships with other measures related to reliability and information theory, see Rajesh and Sunoj (2019) and Toomaj and Atabay (2022).
Entropy, CRE and their related measures consider only the quantitative (i.e. probabilistic) information but in many applied fields it is often required to consider qualitative characteristics or the utility of the random events as well. For example, in a two person game, it is necessary to take into account the various random strategies (quantitative) of the players involved as well as the gain or loss (qualitative) corresponding to the chosen strategy. Noticing the importance of qualitative characteristics of the random events in various applied fields, Belis and Guiasu (1968) introduced the concept of weighted entropy by assigning non-negative weights to each event based on their utility. For the continuous case, weighted differential entropy can be defined as
$H^{w}(X)=-\int_{0}^{+\infty}xf(x)\log f(x)dx,$ (7)
where the factor $x$ is the linear weight function that gives more importance to the larger values of the rv $X$. Motivated from the work of weighted entropy measure, numerous weighted measures are introduced in the literature. Misagh et al. (2011) introduced weighted cumulative residual entropy (WCRE) measure which is defined as
$\xi^{w}(X)=-\int_{0}^{+\infty}x\bar{F}(x)\log\bar{F}(x)dx.$ (8)
Chakraborty and Pradhan (2023) proposed weighted cumulative residual Tsallis entropy (WCRTE) as
$\xi_{\alpha}^{w}(X)=\frac{1}{\alpha-1}\int_{0}^{+\infty}x\left(\bar{F}(x)-\bar{F}^{\alpha}(x)\right)dx,\;0<\alpha\neq 1.$ (9)
As $\alpha\to 1$, it reduces to WCRE. Various properties of WCRTE measure along with its dynamic version and its potential as a risk measure for heavy tailed risk analysis was discussed in Chakraborty and Pradhan (2023). WCRTE is further studied by many authors, see, for example, Kattumannil et al. (2022, 2024) and Zuo and Yin (2025).
Estimations of entropy and related measures were considered in literature by many authors; see, for example, Vasicek (1976), Ebrahimi et al. (1994), Correa (1995), Noughabi (2010) and the references therein. Recently, research on weighted information measures gained quite a lot of attention see, for example, Balakrishnan et al. (2022), Chakraborty (2023), Hashempour et al. (2023), Khammar and Jahanshahi (2018) and the references therin. However, estimations of weighted measures have not been studied to the same extent. The main focus of this article is to study various non-parametric estimators of WCRTE measure. Chakraborty and Pradhan (2023) introduced one non-parametric estimator of WCRTE measure. Here we propose four estimators of WCRTE, study their asymptotic properties and compare their performances. Also, we briefly discuss non-parametric estimations of WCRE measure. We develop two uniformity tests based on an estimator of WCRTE and WCRE measures. The rest of the paper is organised as follows:
We study some important properties of WCRTE measure in Section 2. Non-parametric estimators of WCRTE and WCRE are introduced and their asymptotic properties are investigated in Section 3. A simulation study is performed to compare the performance of the estimators in Section 4. Uniformity tests are discussed in Section 5. Finally, some concluding remarks are made in Section 6.
2 Properties of WCRTE
In this section, we study some properties of WCRTE measure. The following theorem discusses the existence of WCRTE measure.
Theorem 2.1.
For a non-negative continuous rv $X$, $\xi_{\alpha}^{w}(X)$ exists if variance of $X$ exists when $\alpha>1$.
Proof.
From Eq. (9), we have
$\xi_{\alpha}^{w}(X)=\frac{1}{\alpha-1}\left[\frac{E(X^{2})}{2}-\int_{0}^{+\infty}x\bar{F}^{\alpha}(x)dx\right].$
Now,
$\int_{0}^{+\infty}x\bar{F}^{\alpha}(x)dx$ $=$ $\int_{0}^{1}x\bar{F}^{\alpha}(x)dx+\int_{1}^{+\infty}x\bar{F}^{\alpha}(x)dx$
$\leq$ $\int_{0}^{1}xdx+\int_{1}^{+\infty}x\left[\frac{E(X^{p})}{x^{p}}\right]^{\alpha}dx$
$=$ $\frac{1}{2}+\left[E(X^{p})\right]^{\alpha}\int_{1}^{+\infty}x^{1-p\alpha}dx.$
$<$ $\infty,$
if $p>2/\alpha$. Therefore, it is easy to see that, for $\alpha>1$, $\xi_{\alpha}^{w}(X)$ exists when $E(X^{2})$ exists i.e. the variance exists. Hence the proof. ∎
We calculate WCRTE for some popular distributions and present them in Table 1.
Next we provide a lower bound for WCRTE measure in terms of entropy.
Theorem 2.2.
For a non-negative continuous rv $X$, the following inequality holds:
$\xi_{\alpha}^{w}(X)\geq\exp\left[H(X)+E(\log X)+\eta(\alpha)\right],$
where $\eta(\alpha)=\int_{0}^{1}\log\frac{u-u^{\alpha}}{\alpha-1}du.$
Proof.
Using log-sum inequality we can write
$\int_{0}^{+\infty}f(x)\log\frac{f(x)}{\frac{1}{\alpha-1}x(\bar{F}(x)-\bar{F}^{\alpha}(x)}dx\geq\int_{0}^{+\infty}f(x)dx\;\log\frac{\int_{0}^{+\infty}f(x)dx}{\xi_{\alpha}^{w}(X)}=-\log\xi_{\alpha}^{w}(X).$ (10)
Now,
$\int_{0}^{+\infty}f(x)\log\frac{f(x)}{\frac{1}{\alpha-1}x(\bar{F}(x)-\bar{F}^{\alpha}(x)}dx$ $=$ $-H(X)-\int_{0}^{+\infty}f(x)\log\left[\frac{1}{\alpha-1}x(\bar{F}(x)-\bar{F}^{\alpha}(x)\right]dx$ (11)
$=$ $-H(X)-\int_{0}^{+\infty}f(x)\log x\;dx$
$\;\;\;\;-\int_{0}^{+\infty}f(x)\log\frac{\bar{F}(x)-\bar{F}^{\alpha}(x)}{\alpha-1}dx$
$=$ $-H(X)-E(\log X)-\int_{0}^{1}\log\frac{u-u^{\alpha}}{\alpha-1}du.$
Combining Eq. (10) and (11) and after some simplifications, we get the result. ∎
Table 1: WCRTE for some distributions.
Distributions $F(x)$ $\xi_{\alpha}^{w}(X)$
[…]
Exponential $1-e^{-\lambda x};\;x>0,\;\lambda>0$ $\frac{\alpha+1}{\alpha\lambda^{2}}$
Rayleigh $1-e^{-\frac{x^{2}}{2\sigma^{2}}};\;x\geq 0,\;\sigma>0$ $\frac{\sigma^{2}}{\alpha}$
Pareto I $1-\left(\frac{k}{x}\right)^{\delta};\;x>k,\;\delta>0$ $\frac{\delta k^{2}}{(\delta-2)(\delta\alpha-2)}$, $\delta>(<)2,\alpha>(<)1$
Weibull $1-e^{-(\lambda x)^{p}},\;x>0,\;\lambda,\;p>0$ $\frac{\Gamma\left(\frac{2}{p}\right)\left(1-\frac{1}{\alpha^{2/p}}\right)}{p\lambda^{2}(\alpha-1)}$
Now we provide an alternative expression for WCRTE which will be later used to develop a non-parametric estimator.
Theorem 2.3.
For a non-negative continuous rv $X$, the following relation holds:
$\xi_{\alpha}^{w}(X)=\frac{1}{2(\alpha-1)}\int_{0}^{+\infty}x^{2}\left(1-\alpha\bar{F}^{\alpha-1}(x)\right)dF(x).$
Proof.
$\frac{1}{2(\alpha-1)}\int_{0}^{+\infty}x^{2}\left(1-\alpha\bar{F}^{\alpha-1}(x)\right)dF(x)$ $=$ $\frac{1}{\alpha-1}\int_{0}^{+\infty}\left(\int_{0}^{x}vdv\right)\left(1-\alpha\bar{F}^{\alpha-1}(x)\right)dF(x)$
$=$ $\frac{1}{\alpha-1}\int_{0}^{+\infty}\left[\int_{v}^{+\infty}\left(1-\alpha\bar{F}^{\alpha-1}(x)\right)dF(x)\right]vdv$
$=$ $\frac{1}{\alpha-1}\int_{0}^{+\infty}v(\bar{F}(v)-\bar{F}^{\alpha}(v))dv.$
Hence the result follows. ∎
3 Non-parametric Estimation
In this section, we develop some non-parametric estimators for WCRTE and WCRE measures and study their asymptotic properties. Let $X_{1},X_{2},...,X_{n}$ be a random sample of size $n$ drawn from a distribution with cdf $F$ and let $X_{(1)}\leq X_{(2)}\leq...\leq X_{(n)}$ be the corresponding order statistics. Further let $F_{n}$ be the empirical distribution function of $X$. Then
$F_{n}(x)=\frac{\text{number of obs.}\leq x}{n}=\begin{cases}0,\;\;\;\;\mbox{if}\;x<X_{(1)},\\ \frac{i}{n},\;\;\;\mbox{if}\;X_{(i)}\leq x<X_{(i+1)},\;i=1,2,\cdots n-1,\\ 1,\;\;\;\mbox{if}\;x\geq X_{(n)}.\end{cases}$
3.1 Estimation of WCRTE
Using empirical distribution function, Chakraborty and Pradhan (2023) proposed an estimator of WCRTE as
$\hat{\xi}_{\alpha}^{w}(X)$ $=$ $\frac{1}{2(\alpha-1)}\int_{0}^{+\infty}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx$ (12)
$=$ $\frac{1}{2(\alpha-1)}\sum_{i=1}^{n-1}\left(X^{2}_{(i+1)}-X^{2}_{(i)}\right)\left[\left(1-\frac{i}{n}\right)-\left(1-\frac{i}{n}\right)^{\alpha}\right].$
Now we propose two new estimators for WCRTE measure. These estimators are motivated from the works of Vasicek (1976) and Ebrahimi et al. (1994). Note that, using the integral transformation $F(x)=u$, we can express WCRTE in Eq. (9) as
$\xi_{\alpha}^{w}(X)$ $=$ $\frac{1}{\alpha-1}\int_{0}^{1}F^{-1}(u)[1-u-(1-u)^{\alpha}]\;\frac{d}{du}\left(F^{-1}(u)\right)du$ (13)
$=$ $\frac{1}{2(\alpha-1)}\int_{0}^{1}[1-u-(1-u)^{\alpha}]\;\frac{d}{du}\left(F^{-1}(u)\right)^{2}du.$
Vasicek (1976) estimated $\frac{d}{du}\left(F^{-1}(u)\right)$ from the slope of the line joining the points $(F(X_{(i+m)}),X_{(i+m)})$ and $(F(X_{(i-m)}),X_{(i-m)})$. The window size $m$ is an integer and $m<\frac{n}{2}$. So we have
$\frac{d}{du}\left(F^{-1}(u)\right)$ $\approx$ $\frac{X_{(i+m)}-X_{(i-m)}}{F(X_{(i+m)})-F(X_{(i-m)})}$
$=$ $\frac{X_{(i+m)}-X_{(i-m)}}{\frac{i+m}{n}-\frac{i-m}{n}}$
$=$ $\frac{n}{2m}(X_{(i+m)}-X_{(i-m)}).$
Note that, $X_{(i)}=X_{(1)}$ if $i<1$ and $X_{(i)}=X_{(n)}$ if $i>n$. Now we can estimate $\frac{d}{du}\left(F^{-1}(u)\right)^{2}$ as
$\frac{d}{du}\left(F^{-1}(u)\right)^{2}$ $=$ $2F^{-1}(u)\frac{d}{du}\left(F^{-1}(u)\right)$
$\approx$ $\frac{n}{2m}(X_{(i+m)}+X_{(i-m)})(X_{(i+m)}-X_{(i-m)})$
[…]
Therefore, a Vasicek-type estimator for WCRTE is defined as
[…]
$=$ $\frac{1}{4m(\alpha-1)}\sum_{i=1}^{n}\left(X^{2}_{(i+m)}-X^{2}_{(i-m)}\right)\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right].$
Here $V$ in $\xi^{w}_{\alpha}(V)$ stands to emphasize that this estimator is a Vasicek-type estimator.
Ebrahimi et al. (1994) modified Vasicek entropy estimator with the help of a weight function in order to estimate the slope more accurately at the extreme points. The weight function of Ebrahimi is
$C_{i}=\begin{cases}1+\frac{i-1}{m},\;\;\;\;\mbox{if}\;\;\;1\leq i\leq m,\\ 2,\;\;\;\;\;\;\;\;\;\;\;\;\;\mbox{if}\;\;\;m+1\leq i\leq n-m,\\ 1+\frac{n-i}{m},\;\;\;\;\mbox{if}\;\;\;n-m+1\leq i\leq n.\end{cases}$
Using these weights, another estimator of WCRTE can be defined as
[…]
$=$ $\frac{1}{2m(\alpha-1)}\sum_{i=1}^{n}\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C_{i}}\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right].$
Here $E$ stands for Ebrahimi as $V$ stands for Vasicek in (14).
We introduce another estimator of WCRTE measure by modifying Ebrahimi weights in $\xi_{\alpha}^{w}E$. The modifications happen in the estimation of $\frac{d}{du}\left(F^{-1}(u)\right)^{2}$ as follows:
$\frac{d}{du}\left(F^{-1}(u)\right)^{2}$ $=$ $2F^{-1}(u)\frac{d}{du}\left(F^{-1}(u)\right)$
$\approx$ $2\frac{(X_{(i+m)}+X_{(i-m)})}{C_{i}}\frac{n}{C_{i}m}(X_{(i+m)}-X_{(i-m)})$
$=$ $\frac{2n}{m}\left(\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C^{2}_{i}}\right).$
Here the modifications are made in the estimations of $\frac{d}{du}\left(F^{-1}(u)\right)^{2}$ as well as $F^{-1}(u)$. The new estimator is given by
$\xi_{\alpha}^{w}N$ $=$ $\frac{1}{2(\alpha-1)}\frac{1}{n}\sum_{i=1}^{n}\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right]\frac{2n}{m}\left(\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C^{2}_{i}}\right)$ (16)
$=$ $\frac{1}{m(\alpha-1)}\sum_{i=1}^{n}\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C^{2}_{i}}\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right].$
The following theorem shows the consistency of these estimators. Proof is ovbious from the works of Vasicek (1976), and hence omitted.
Theorem 3.1.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample from a continuous distribution having cdf $F$, pdf $f$ and finite second moment. Then, for $\alpha>1$,
  1. 1.
    $\xi_{\alpha}^{w}V\stackrel{{\scriptstyle p}}{{\to}}\xi_{\alpha}^{w}(X)$,
  2. 2.
    $\xi_{\alpha}^{w}E\stackrel{{\scriptstyle p}}{{\to}}\xi_{\alpha}^{w}(X)$,
  3. 3.
    $\xi_{\alpha}^{w}N\stackrel{{\scriptstyle p}}{{\to}}\xi_{\alpha}^{w}(X)$,
as $m,n\to+\infty$, such that $\frac{m}{n}\to 0$. The notation $\stackrel{{\scriptstyle p}}{{\to}}$” indicates the convergence in probability.
Next, we define an estimator for WCRTE, by making use of Theorem 2.3, as
$\xi_{\alpha}^{w}L$ $=$ $\frac{1}{2(\alpha-1)}\int_{0}^{+\infty}x^{2}\left(1-\alpha\bar{F}_{n}^{\alpha-1}(x)\right)dF_{n}(x)$ (17)
$=$ $\frac{1}{2(\alpha-1)}\frac{1}{n}\sum_{i=1}^{n}X^{2}_{(i)}\left[1-\alpha\left(1-\frac{i}{n}\right)^{\alpha-1}\right].$
This estimator belongs to the class of estimators called linear combination of functions of order statistics.
The following proposition discusses the effect of scale transformation on the estimators for WCRTE measure.
Proposition 3.1.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample with WCRTE $\xi_{\alpha}^{w}(X)$ and let $Y_{i}=\theta X_{i},$ $\theta>0,\;i=1(1)n$. Suppose the estimators for WCRTE of $X$ and $Y$ are $\xi_{\alpha}^{w}V(X),\;\xi_{\alpha}^{w}E(X),\;\xi_{\alpha}^{w}N(X),\;\xi_{\alpha}^{w}L(X)$ and $\xi_{\alpha}^{w}V(Y),\;\xi_{\alpha}^{w}E(Y),\;\xi_{\alpha}^{w}N(Y),\;\xi_{\alpha}^{w}L(Y)$, respectively. Then
  1. 1.
    $\xi_{\alpha}^{w}V(Y)=\theta^{2}\xi_{\alpha}^{w}V(X)$;
  2. 2.
    $\xi_{\alpha}^{w}E(Y)=\theta^{2}\xi_{\alpha}^{w}E(X)$;
  3. 3.
    $\xi_{\alpha}^{w}N(Y)=\theta^{2}\xi_{\alpha}^{w}N(X)$;
  4. 4.
    $\xi_{\alpha}^{w}L(Y)=\theta^{2}\xi_{\alpha}^{w}L(X)$.
Proof.
From Eq. (14), we have
$\xi_{\alpha}^{w}V(Y)$ $=$ $\frac{1}{4m(\alpha-1)}\sum_{i=1}^{n}\left(Y^{2}_{(i+m)}-Y^{2}_{(i-m)}\right)\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right]$
$=$ $\frac{1}{4m(\alpha-1)}\sum_{i=1}^{n}\left(\theta^{2}X_{(i+m)}-\theta^{2}X_{(i-m)}\right)\left[1-\frac{i}{n}-\left(1-\frac{i}{n}\right)^{\alpha}\right]$
$=$ $\theta^{2}\xi_{\alpha}^{w}V(X).$
Similarly, we can prove the other results. ∎
3.1.1 Asymptotic normality of $\xi^{w}_{\alpha}L$
Here we study the asymptotic normality of $\xi_{\alpha}L$. Consider
$T_{n}=2(\alpha-1)\xi_{\alpha}^{w}L$ $=$ $\frac{1}{n}\sum_{i=1}^{n}X^{2}_{(i)}\left[1-\alpha\left(1-\frac{i}{n+1}\right)^{\alpha-1}\right]$ (18)
$=$ $\frac{1}{n}\sum_{i=1}^{n}\psi\left(\frac{i}{n+1}\right)h(X_{(i)}),$
where $\psi(u)=1-\alpha(1-u)^{\alpha-1}$ and $h(u)=u^{2}$. The following theorem discusses the asymptotic normality of $T_{n}$ and hence of $\xi_{\alpha}^{w}L$. The notation $\stackrel{{\scriptstyle d}}{{\to}}$” indicates convergence in distribution.
Theorem 3.2.
For a non-negative, absolutely continuous rv $X$, $\sqrt{n}\left(\xi_{\alpha}^{w}L-\xi_{\alpha}^{w}(X)\right)$ is asymptotically normally distributed with zero mean and variance
$\sigma^{2}=\frac{2}{(\alpha-1)^{2}}\int_{0}^{+\infty}\int_{0}^{+\infty}xyF(x)\bar{F}(y)\left[1-\alpha\bar{F}^{\alpha-1}(x)\right]\left[1-\alpha\bar{F}^{\alpha-1}(y)\right]dxdy.$
Proof.
From the results given in Chernoff et al. (1967) and Shorack (1969), it can be shown that $T_{n}$ converges to normal distribution since $\sqrt{n}(T_{n}-\mu)\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma_{1}^{2})$ as $n\to+\infty$. Note that, $E(T_{n})=\mu$ and it can be obtained as
$\mu$ $=$ $\int_{0}^{1}\psi(u)h(F^{-1}(u))du$ (19)
$=$ $\int_{0}^{1}\left(1-\alpha(1-u)^{\alpha-1}\right)(F^{-1}(u))^{2}du$
$=$ $2(\alpha-1)\xi_{\alpha}^{w}(X).$
The variance of $T_{n}$ is
$\sigma_{1}^{2}$ $=$ $2\int_{0}^{1}\int_{u}^{1}\psi(u)\psi(w)\left(h(F^{-1}(u))\right)^{\prime}\left(h(F^{-1}(w))\right)^{\prime}u(1-w)dwdu$ (20)
$=$ $8\int_{0}^{1}\int_{u}^{1}\psi(u)\psi(w)u(1-w)F^{-1}(u)F^{-1}(w)dF^{-1}(w)dF^{-1}(u)$
$=$ $8\int_{0}^{+\infty}\int_{y}^{+\infty}xyF(y)\bar{F}(x)\left[1-\alpha\bar{F}^{\alpha-1}(x)\right]\left[1-\alpha\bar{F}^{\alpha-1}(y)\right]dxdy.$
The result follows from equations (19) and (20) and by using the property of normal distribution. ∎
Since $\sigma^{2}$ contains unknown parameters in $F$, one needs to work with its estimator for pracical considerations. A consistent estimator for $\sigma^{2}$ based on a data set can be defined as
$\hat{\sigma}^{2}$ $=$ $\frac{2}{(\alpha-1)^{2}}\sum_{j=1}^{n-1}\sum_{i=j+1}^{n-1}\frac{j}{n}\left(1-\frac{i}{n}\right)\left[1-\alpha\left(1-\frac{i}{n}\right)^{\alpha-1}\right]\left[1-\alpha\left(1-\frac{j}{n}\right)^{\alpha-1}\right]$
[…]
3.2 Estimation of WCRE
WCRE is a limiting case of WCRTE measure. Like we expresssed WCRTE in Eq. (13) by using integral transformation $F(x)=u$, we can similarly expressed WCRE as
$\xi^{w}(X)=-\frac{1}{2}\int_{0}^{1}(1-u)\log(1-u)\;\frac{d}{du}\left(F^{-1}(u)\right)^{2}du.$ (21)
From Eq. (21), we can develop the estimators for WCRE. The Vasicek-type and the Ebrahimi-type estimators of WCRE can be obtained as
$\xi^{w}V=-\frac{1}{4m}\sum_{i=1}^{n-1}\left(X^{2}_{(i+m)}-X^{2}_{(i-m)}\right)\left(1-\frac{i}{n}\right)\log\left(1-\frac{i}{n}\right);$
and
$\xi^{w}E=-\frac{1}{2m}\sum_{i=1}^{n-1}\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C_{i}}\left(1-\frac{i}{n}\right)\log\left(1-\frac{i}{n}\right),$
respectively. We define a new estimator for WCRE as
$\xi^{w}N=-\frac{1}{m}\sum_{i=1}^{n-1}\frac{X^{2}_{(i+m)}-X^{2}_{(i-m)}}{C^{2}_{i}}\left(1-\frac{i}{n}\right)\log\left(1-\frac{i}{n}\right).$
Remark 3.1.
As $\alpha\to 1$, $\xi_{\alpha}^{w}(X)\to\xi^{w}(X)$. This limiting condition also holds for the respective estimators as well. To be specific, $\xi_{\alpha}^{w}V\to\xi^{w}V$, $\xi_{\alpha}^{w}E\to\xi^{w}E$ and $\xi_{\alpha}^{w}N\to\xi^{w}N$ as $\alpha\to 1$.
Theorem 3.3.
For a non-negative continuous rv $X$,
$\xi^{w}(X)=-\frac{1}{2}\int_{0}^{+\infty}x^{2}\left(1+\log\bar{F}(x)\right)dF(x.)$
Proof.
Proof follows by proceeding along the same line as in Theorem 2.3. ∎
Using Theorem 3.3, we define an estimator for WCRE as
$\xi^{w}L$ $=$ $-\frac{1}{2n}\sum_{i=1}^{n}X^{2}_{(i)}\left[1+\log\left(1-\frac{i}{n}\right)\right].$ (22)
In the following proposition, we discuss the effect of scale transformation on the estimators of WCRE measure. Proof is similar to that of Proposition 3.1, and hence omitted.
Proposition 3.2.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample with WCRE $\xi^{w}(X)$ and let $Y_{i}=\theta X_{i},\;\theta>0,\;i=1(1)n$. Suppose the estimators for WCRE of $X$ and $Y$ are $\xi^{w}V(X)$, $\xi^{w}E(X)$, $\xi^{w}N(X)$, $\xi^{w}L(X)$ and $\xi^{w}V(Y),\;\xi^{w}E(Y),\;\xi^{w}N(Y),\;\xi^{w}L(Y)$, respectively. Then
  1. (i)
    $\xi^{w}V(Y)=\theta^{2}\xi^{w}V(X)$;
  2. (ii)
    $\xi^{w}E(Y)=\theta^{2}\xi^{w}E(X)$;
  3. (iii)
    $\xi^{w}N(Y)=\theta^{2}\xi^{w}N(X)$;
  4. (iv)
    $\xi^{w}L(Y)=\theta^{2}\xi^{w}L(X)$.
In the next theorem, we discuss the asymptotic normality of the estimator $\xi^{w}L$.
Theorem 3.4.
For a non-negative continuous rv $X$, $\sqrt{n}\left(\xi^{w}L-\xi^{w}(X)\right)$ converges to $N(0,\sigma^{2})$ as $n\to+\infty$ where
$\sigma^{2}=2\int_{0}^{+\infty}\int_{y}^{+\infty}xyF(y)\bar{F}(x)\left[1+\log\bar{F}(x)\right]\left[1+\log\bar{F}(y)\right]dxdy.$
Proof.
Consider, $\zeta_{n}=-2\xi^{w}L$. Now proceeding alomg the same lines as in Theorem 3.2, it can be shown that $\sqrt{n}(\zeta_{n}-\mu)\to N(0,\sigma_{1}^{2})$ as $n\to+\infty$, where the mean is $\mu=-2\xi^{w}(X)$ and the variance is
$\sigma_{1}^{2}=8\int_{0}^{+\infty}\int_{y}^{+\infty}xyF(y)\bar{F}(x)\left[1+\log\bar{F}(x)\right]\left[1+\log\bar{F}(y)\right]dxdy.$
Hence the proof follows. ∎
A consistent estimator for $\sigma^{2}$ can be obtained as
$\hat{\sigma^{2}}$ $=$ $2\sum_{j=1}^{n-1}\sum_{i=j+1}^{n-1}\frac{j}{n}\left(1-\frac{i}{n}\right)\left(1+\log\left(1-\frac{i}{n}\right)\right)\left(1+\log\left(1-\frac{j}{n}\right)\right)$
[…]
4 Comparison of estimators
We conduct a simulation study to assess the performance of the proposed estimators. We compare the performance of these estimators by means of Bias and mean square error (MSE). For reference distributions, we consider exponential, Weibull and uniform distributions. We denote them by Exp($\lambda$), WE($\alpha,\beta$) and U($a,b$), respectively. The cdfs of these distributions are provided in Table 1. We generate 10000 random samples of size $n$ = 10, 20 and 30 from Exp(1), Exp(2), U(0,1) and WE(2,1) distributions and calculate bias and MSEs of these estimators. The results are provided in Tables 2-6.
Table 2: Bias and MSE of $\hat{\xi}_{\alpha}^{w}(X)$ and $\xi_{\alpha}^{w}L$ for $\alpha$ = 2.
$\hat{\xi}_{\alpha}^{w}(X)$ $\xi_{\alpha}^{w}L$
Distributions $n$ Bias MSE Bias MSE
10 -0.8259 0.9686 -0.7251 0.8848
Exp(1) 20 -0.7879 0.7751 -0.7400 0.7205
30 -0.7749 0.7143 -0.7427 0.6693
10 -0.2076 0.0616 -0.1814 0.0560
Exp(2) 20 -0.1964 0.0491 -0.1836 0.0451
30 -0.1940 0.0443 -0.1847 0.0416
10 -0.0086 0.00032 -0.0086 0.00043
U(0,1) 20 -0.0043 0.00014 0.0041 0.00017
30 -0.0029 0.00009 0.0028 0.00010
10 -0.0250 0.0078 0.0251 0.0102
WE(2,1) 20 -0.0120 0.0040 0.0116 0.0046
30 -0.0081 0.0027 0.0074 0.0030
Table 3: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for Exp(1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -1.005 1.1386 -1.003 1.1301 -1.001 1.1314
2 -0.9764 1.1045 -0.9301 1.0660 -0.8945 1.0177
3 -0.9564 1.1038 -0.8787 1.0321 -0.7865 0.9478
4 -0.9365 1.0956 -0.8373 0.9963 -0.6966 0.9469
20 1 -0.9313 0.9486 -0.9315 0.9568 -0.9281 0.9495
2 -0.8965 0.9092 -0.8801 0.8866 -0.8489 0.8500
3 -0.8653 0.8713 -0.8234 0.8213 -0.7680 0.7732
4 -0.8379 0.8515 -0.7746 0.7817 -0.6852 0.7200
5 -0.8146 0.8290 -0.7223 0.7490 -0.6122 0.6816
6 -0.7804 0.8134 -0.6824 0.7354 -0.5342 0.7161
7 -0.7609 0.8166 -0.6364 0.7442 -0.4705 0.7278
8 -0.7490 0.8090 -0.5964 0.7315 -0.4171 0.6940
9 -0.7405 0.8078 -0.5893 0.7189 -0.3539 0.7440
30 1 -0.8943 0.8679 -0.8910 0.8591 -0.8951 0.8664
2 -0.8701 0.8334 -0.8503 0.8087 -0.8287 0.7801
3 -0.8420 0.7972 -0.8062 0.7529 -0.7595 0.7047
4 -0.8177 0.7709 -0.7566 0.7080 -0.6917 0.6440
5 -0.7860 0.7406 -0.7174 0.6706 -0.6287 0.6050
6 -0.7522 0.7716 -0.6673 0.6435 -0.5634 0.5837
7 -0.7283 0.7019 -0.6328 0.6234 -0.4957 0.5722
8 -0.7120 0.6864 -0.5855 0.6152 -0.4329 0.5796
9 -0.6765 0.6700 -0.5413 0.6180 -0.3744 0.5818
10 -0.6584 0.6712 -0.5212 0.5994 -0.3284 0.6004
11 -0.6430 0.6522 -0.4890 0.5917 -0.2638 0.6726
12 -0.6272 0.6498 -0.4635 0.6047 -0.2185 0.7171
13 -0.6116 0.6780 -0.4179 0.6285 -0.1679 0.7245
14 -0.5919 0.6682 -0.4083 0.6152 -0.1098 0.7953
The parameter of WCRTE is taken as $\alpha$ = 2. We choose $\alpha$ = 2 because from Theorem 2.1 we observe that, for a continuous distribution, WCRTE exists if variance of that distribution exists when $\alpha>$1. When $\alpha<$1, we will need higher order moments to exist for WCRTE to exist. In practice, it is better to work with $\alpha>$1, since higher the value of $\alpha$, it is more likely for WCRTE to exist (see Theorem 2.1).
Table 4: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for Exp(2) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.2504 0.0713 -0.2497 0.0706 -0.2506 0.0707
2 -0.2444 0.0697 -0.2366 0.0671 -0.2206 0.0630
3 -0.2371 0.0684 -0.2186 0.0646 -0.1955 0.0595
4 -0.2347 0.0682 -0.2097 0.0640 -0.1734 0.0596
20 1 -0.2327 0.0597 -0.2323 0.0593 -0.2312 0.0589
2 -0.2238 0.0566 -0.2192 0.0550 -0.2103 0.0526
3 -0.2172 0.0553 -0.2054 0.0515 -0.1917 0.0486
4 -0.2097 0.0531 -0.1935 0.0493 -0.1712 0.0450
5 -0.2026 0.0524 -0.1796 0.0473 -0.1529 0.0445
6 -0.1962 0.0513 -0.1684 0.0466 -0.1348 0.0442
7 -0.1912 0.0508 -0.1606 0.0454 -0.1158 0.0454
8 -0.1868 0.0507 -0.1514 0.0458 -0.0985 0.0447
9 -0.1837 0.0509 -0.1450 0.0455 -0.0864 0.0480
30 1 -0.2234 0.0541 -0.2242 0.0544 -0.2247 0.0546
2 -0.2176 0.0520 -0.2129 0.0504 -0.2078 0.0489
3 -0.2084 0.0494 -0.2038 0.0478 -0.1904 0.0443
4 -0.2025 0.0474 -0.1904 0.0443 -0.1748 0.0408
5 -0.1960 0.0459 -0.1794 0.0422 -0.1549 0.0378
6 -0.1883 0.0441 -0.1672 0.0400 -0.1408 0.0368
7 -0.1830 0.0433 -0.1573 0.0386 -0.1237 0.0361
8 -0.1753 0.0421 -0.1486 0.0377 -0.1078 0.0361
9 -0.1716 0.0423 -0.1381 0.0372 -0.0944 0.0382
10 -0.1642 0.0414 -0.1293 0.0378 -0.0776 0.0394
11 -0.1610 0.0412 -0.1209 0.0391 -0.0671 0.0401
12 -0.1576 0.0411 -0.1156 0.0389 -0.0565 0.0418
13 -0.1537 0.0412 -0.1022 0.0394 -0.0404 0.0440
14 -0.1493 0.0414 -0.1008 0.0393 -0.0306 0.0487
In Table 2, we present the bias and the MSEs of $\hat{\xi}^{w}_{\alpha}(X)$ and $\xi_{\alpha}^{w}L$ for these reference distributions. These two estimators do not require a window size. From Table 2, we observe that $\xi_{\alpha}^{w}L$ performs better than $\hat{\xi}^{w}_{\alpha}(X)$ for Exp(1) and Exp(2) distribution. But $\hat{\xi}^{w}_{\alpha}(X)$ performs better than $\xi_{\alpha}^{w}L$ for other two models. However, the difference is negligible for $n\geq 20$. Also, as sample size increases bias and MSE decrease.
The estimators $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ depend on the window size $m(<\frac{n}{2})$. We consider all possible values of $m$ for bias and MSE calculation and try to provide a guideline for the choice of $m$. The choice of $m$ will depend on the underlying distribution and the sample size $n$. We highlight the minimum values of MSEs of these estimators in Tables 3-6.
Table 5: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for U(0,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.0152 0.00050 -0.0145 0.00048 -0.0127 0.00043
2 -0.0171 0.00052 -0.0137 0.00043 -0.0086 0.00034
3 -0.0200 0.00058 -0.0129 0.00039 -0.0025 0.00028
4 -0.0236 0.00070 -0.0119 0.00035 0.0047 0.00034
20 1 -0.0077 0.00019 -0.0078 0.00020 -0.0076 0.00019
2 -0.0085 0.00020 -0.0077 0.00018 -0.0065 0.00017
3 -0.0090 0.00020 -0.0075 0.00017 -0.0049 0.00014
4 -0.0104 0.00022 -0.0073 0.00016 -0.0029 0.00011
5 -0.0118 0.00023 -0.0070 0.00015 -0.00035 0.00010
6 -0.0135 0.00027 -0.0070 0.00014 0.0026 0.00010
7 -0.0154 0.00031 -0.0067 0.00013 0.0056 0.00014
8 -0.0175 0.00036 -0.0064 0.00012 0.0092 0.00020
9 -0.0197 0.00044 -0.0061 0.00011 0.0130 0.00029
30 1 -0.0053 0.00011 -0.0052 0.00011 -0.0051 0.00011
2 -0.0055 0.00011 -0.0052 0.00011 -0.0046 0.00010
3 -0.0061 0.00012 -0.0051 0.00011 -0.0041 0.00009
4 -0.0066 0.00012 -0.0051 0.00010 -0.0031 0.00008
5 -0.0071 0.00012 -0.0050 0.00009 -0.0020 0.00007
6 -0.0081 0.00013 -0.0048 0.00008 -0.0007 0.00006
7 -0.0089 0.00014 -0.0050 0.00008 0.0008 0.00006
8 -0.0100 0.00015 -0.0048 0.00008 0.0025 0.00006
9 -0.0112 0.00017 -0.0048 0.00008 0.0045 0.00008
10 -0.0126 0.00020 -0.0046 0.00007 0.0065 0.00010
11 -0.0140 0.00023 -0.0045 0.00007 0.0086 0.00013
12 -0.0154 0.00027 -0.0043 0.00006 0.0110 0.00018
13 -0.0167 0.00031 -0.0041 0.00006 0.0133 0.00024
14 -0.0184 0.00036 -0.0043 0.00006 0.0160 0.00032
From Tables 3 and 4, it is observed that, for Exp(1) and Exp(2) distributions, $\xi_{\alpha}^{w}N$ performs better than other two estimators and $\xi_{\alpha}^{w}V$ performs the worst. All these estimators perform much better for Exp(2) distribution than Exp(1). For Exp(2) distribution, the difference between the MSEs of these estimators is negligible. The optimal choice of $m$ (for different $n$) for each estimator can be obtained from the respective tables. It is not possible to provide an exact formula for the choice of $m$ but approximate formulae may be provided from Tables 3-4, which will provide optimal or near optimal choices of $m$ for exponential distributions. Note from the tables that
  • (a)
    for $\xi_{\alpha}^{w}V$ and $\xi_{\alpha}^{w}E$, when $n\leq$ 20, choose $m=[\frac{n}{2}]-1$. When $n$ = 30, choose $m=[\frac{n}{3}]$;
  • (b)
    for $\xi_{\alpha}^{w}N$, choose $m=[\frac{n}{4}]+1$, for all $n$.
From Table 5 it is observed that, for U(0,1) distribution $\xi_{\alpha}^{w}N$ performs the best and $\xi_{\alpha}^{w}V$ performs the worst, although the performance of these estimators is similar. Performance of these estimators do not significantly vary for the choices of $m$ as well. Also, for U(0,1), these estimators perform better than all the other models considered. Here small values of $m$ will be appropriate.
Table 6: Bias and MSE of $\xi_{\alpha}^{w}V$, $\xi_{\alpha}^{w}E$ and $\xi_{\alpha}^{w}N$ for WE(2,1) distribution with $\alpha$ = 2, $n$ = 10, 20 and 30 and for various $m$.
$\xi_{\alpha}^{w}V$ $\xi_{\alpha}^{w}E$ $\xi_{\alpha}^{w}N$
$n$ $m$ Bias MSE Bias MSE Bias MSE
10 1 -0.0582 0.0081 -0.0562 0.0078 -0.0490 0.0072
2 -0.0588 0.0081 -0.0452 0.0073 -0.0274 0.0068
3 -0.0612 0.0086 -0.0381 0.0074 -0.0038 0.0080
4 -0.0672 0.0092 -0.0305 0.0077 0.0207 0.0107
20 1 -0.0356 0.0042 -0.0347 0.0042 -0.0327 0.0040
2 -0.0328 0.0042 -0.0277 0.0039 -0.0210 0.0038
3 -0.0310 0.0041 -0.0206 0.0039 -0.0069 0.0040
4 -0.0295 0.0041 -0.0145 0.0041 -0.0062 0.0047
5 -0.0296 0.0043 -0.0079 0.0043 -0.0195 0.0059
6 -0.0285 0.0043 -0.0023 0.0047 -0.0343 0.0074
7 -0.0300 0.0047 -0.0030 0.0052 -0.0490 0.0097
8 -0.0310 0.0049 0.0078 0.0055 -0.0634 0.0126
9 -0.0328 0.0052 0.0120 0.0062 -0.0780 0.0154
30 1 -0.0268 0.0029 -0.0260 0.0028 -0.0251 0.0028
2 -0.0234 0.0027 -0.0207 0.0027 -0.0165 0.0026
3 -0.0235 0.0028 -0.0164 0.0027 -0.0080 0.0027
4 -0.0194 0.0027 -0.0099 0.0027 0.0035 0.0031
5 -0.0174 0.0027 -0.0044 0.0030 0.0130 0.0038
6 -0.0155 0.0028 0.0009 0.0031 0.0240 0.0045
7 -0.0139 0.0030 0.0064 0.0035 0.0345 0.0058
8 -0.0177 0.0032 0.0114 0.0040 0.0453 0.0073
9 -0.0119 0.0033 0.0148 0.0043 0.0577 0.0092
10 -0.0114 0.0034 0.0210 0.0049 0.0676 0.0111
11 -0.0122 0.0036 0.0248 0.0054 0.0780 0.0132
12 -0.0117 0.0037 0.0288 0.0057 0.0877 0.0154
13 -0.0138 0.0038 0.0335 0.0066 0.0985 0.0180
14 -0.0149 0.0040 0.0359 0.0069 0.1104 0.0213
From Table 6, we observe that for WE(2,1) (Rayleigh) distribution, all estimators perform similarly, although the performance of $\xi_{\alpha}^{w}N$ may be considered to be the best whereas $\xi_{\alpha}^{w}V$ is the worst. However the difference in performance may be considered to be insignificant. Thus we prescribe to choose small $n$ for better result.
From the simulation study we find that $\xi_{\alpha}^{w}N$ performs the best, $\xi_{\alpha}^{w}V$ performs the worst and performance of $\xi_{\alpha}^{w}E$ stays in the middle for all the distributions considered. Also they perform better than $\hat{\xi}_{\alpha}^{w}(X)$ and $\xi_{\alpha}^{w}L$ in most of the cases. When sample size increases MSEs of all estimators decrease which is obvious. In general, $\xi_{\alpha}^{w}N$ has the lowest MSE among all other estimators. However, $\xi_{\alpha}^{w}N$ is more sensitive to the choice of $m$ than $\xi_{\alpha}^{w}V$ and $\xi_{\alpha}^{w}E$, whereas $\xi_{\alpha}^{w}V$ is the least sensitive to the choice of $m$. So, in order to work with $\xi_{\alpha}^{w}N$ and $\xi_{\alpha}^{w}E$, appropriate choice of $m$ is important for better results. Also, if the sample size is large then all estimators perform similarly. So, it is better to use $\xi_{\alpha}^{w}L$ for large $n$ because its asymptotic distribution is normal. This will be very useful in various statistical inference problems.
5 Applications
In this section, we develop a goodness-of-fit test for uniform distribution using an estimator of WCRTE measure. Uniform distribution is used for random number generation. Uniformity test is an important problem in statistics and this problem has been addressed by many researchers in the literature. For example, Stephens (1974) studied uniformity test using cdf-based statistic and Dudewicz and Van Der Meulen (1981) proposed an entropy-based test of uniformity. For recent development in this area, redears may refer to Noughabi (2022), Chakraborty et al. (2023) and the references therin.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample from a continuous distribution with pdf $f$ concentrated on [0,1], i.e. $f(x)=0$ if $x\notin$[0,1]. Suppose $X_{(1)},X_{(2)},\cdots,X_{(n)}$ are the corresponding order statistics. We want to test whether the data come from a U(0,1) distribution. So we want to test
$H_{0}:f(x)\sim U(0,1)\;\;\;\;vs.\;\;\;\;H_{1}:f(x)\not\sim U(0,1).$
Consider the estimator of WCRTE
$\hat{\xi}_{\alpha}^{w}(X)=\frac{1}{2(\alpha-1)}\sum_{i=1}^{n-1}\left(X^{2}_{(i+1)}-X^{2}_{(i)}\right)\left[\left(1-\frac{i}{n}\right)-\left(1-\frac{i}{n}\right)^{\alpha}\right].$
This estimator was introduced by Chakraborty and Pradhan (2023). We will use this estimator as a test statistic. Note that, when $\alpha\to 1$, $\hat{\xi}_{\alpha}^{w}(X)$ reduces to $\hat{\xi}^{w}(X)$ where $\hat{\xi}^{w}(X)$ is the estimator of WCRE (Misagh et al., 2011).
The following lemma will be useful for the construction of the test.
Lemma 5.1.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample drawn from a continuous distribution concentrated on [0,1]. Then, $0\leq\hat{\xi}_{\alpha}^{w}(X)\leq\frac{1}{2\alpha^{\frac{\alpha}{\alpha-1}}}$ for $\alpha>1$.
Proof.
Note that, $f(v)=(1-v)-(1-v)^{\alpha},\,0<v<1$, has the maximum value $\frac{\alpha-1}{\alpha^{\frac{\alpha}{\alpha-1}}}$ for $\alpha>1$. So we have,
$\hat{\xi}^{w}_{\alpha}(X)\leq\frac{1}{2(\alpha-1)}\frac{(\alpha-1)}{\alpha^{\frac{\alpha}{\alpha-1}}}\sum_{i=1}^{n-1}(X^{2}_{(i+1)}-X^{2}_{(i)})=\frac{1}{2\alpha^{\frac{\alpha}{\alpha-1}}}(X^{2}_{(n)}-X^{2}_{(1)})\leq\frac{1}{2\alpha^{\frac{\alpha}{\alpha-1}}}.$
Hence the result follows. ∎
Under $H_{0}$, $\xi_{\alpha}^{w}(X)=\frac{(\alpha+4)}{6(\alpha+1)(\alpha+2)}$, which lies between $\left(0,\frac{1}{2\alpha^{\frac{\alpha}{\alpha-1}}}\right)$. Therefore, we can use $\hat{\xi}_{\alpha}^{w}(X)$ as a test statistic and we will reject the null hypothesis when $\hat{\xi}_{\alpha}^{w}(X)$ is large or small.The critical region for a sample of size $n$ and significance level $\gamma$ is
$\hat{\xi}_{\alpha}^{w}(X)\leq\mathcal{C}_{\frac{\gamma}{2},n}\;\;or\;\;\hat{\xi}_{\alpha}^{w}(X)\geq\mathcal{C}_{1-\frac{\gamma}{2},n},$
where $\mathcal{C}_{\gamma,n}(\mathcal{C}_{1-\gamma,n})$ is the lower (upper) quantile point of the empirical distribution of $\hat{\xi}_{\alpha}^{w}(X)$. Note that, when $\alpha=6.586506$, the value of $\xi_{\alpha}^{w}(X)$ under $H_{0}$ lies in the middle of $\left(0,\frac{1}{2\alpha^{\frac{\alpha}{\alpha-1}}}\right)$.
Table 7: Critical values of $\hat{\xi}^{w}_{\alpha}(X)$.
$n$ $\alpha$ $\mathcal{C}_{0.025,n}$ $\mathcal{C}_{0.975,n}$ $n$ $\alpha$ $\mathcal{C}_{0.05,n}$ $\mathcal{C}_{0.95,n}$
10 1 0.06755 0.15720 15 1 0.08561 0.15572
2 0.04104 0.10149 2 0.05006 0.09995
5 0.01700 0.04822 5 0.02056 0.04663
7 0.01175 0.03553 7 0.01431 0.03407
10 0.00823 0.02545 10 0.00992 0.02442
20 1 0.09670 0.15478 25 1 0.10236 0.15404
2 0.05553 0.09875 2 0.05888 0.09742
5 0.02248 0.04581 5 0.02383 0.04495
7 0.01584 0.03327 7 0.01691 0.03273
10 0.01093 0.02360 10 0.01170 0.02300
30 1 0.10791 0.15291 35 1 0.10948 0.15209
2 0.06160 0.09629 2 0.06362 0.09556
5 0.02508 0.04392 5 0.02594 0.04355
7 0.01780 0.03209 7 0.01840 0.03158
10 0.01215 0.02252 10 0.01265 0.02226
40 1 0.11240 0.15153 45 1 0.11442 0.15112
2 0.06460 0.09505 2 0.06611 0.09450
5 0.02643 0.04316 5 0.02730 0.04286
7 0.01883 0.03137 7 0.01925 0.03116
10 0.01294 0.02202 10 0.01320 0.02180
50 1 0.11610 0.15061 60 1 0.11894 0.15022
2 0.06712 0.09420 2 0.06870 0.09370
5 0.02758 0.04248 5 0.02850 0.04196
7 0.01940 0.03073 7 0.02021 0.03044
10 0.01347 0.02167 10 0.01392 0.02121
70 1 0.12029 0.14941 80 1 0.12218 0.14885
2 0.06980 0.09252 2 0.07075 0.09232
5 0.02884 0.04162 5 0.02927 0.04134
7 0.02051 0.03004 7 0.02073 0.02971
10 0.01416 0.02106 10 0.01429 0.02082
90 1 0.12325 0.14842 100 1 0.12441 0.14791
2 0.07165 0.09202 2 0.07242 0.09160
5 0.02974 0.04096 5 0.02999 0.04071
7 0.02103 0.02955 7 0.02119 0.02926
10 0.01448 0.02064 10 0.01474 0.02050
Lemma 5.2.
The test based on $\hat{\xi}_{\alpha}^{w}(X)$ is consistent.
Proof.
We have,
$\hat{\xi}_{\alpha}^{w}(X)$ $=$ $\frac{1}{\alpha-1}\int_{0}^{+\infty}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx$ (23)
$=$ $\frac{1}{\alpha-1}\left[\int_{0}^{1}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx+\int_{1}^{+\infty}x\left(\bar{F}_{n}(x)-\bar{F}^{\alpha}_{n}(x)\right)dx\right].$
Using Glivenco-Cantelli theorem, the first integral in Eq. (23) converges to
$\int_{0}^{1}x(\bar{F}(x)-\bar{F}^{\alpha}(x))dx$.
We have from Rao et al. (2004), for $x\in[1,+\infty)$,
$\bar{F}_{n}(x)\leq x^{-p}\left(\sup_{n}\frac{1}{n}\sum_{i=1}^{n}X_{i}^{p}\right).$
So, $\int_{1}^{+\infty}x\bar{F}_{n}(x)dx\leq\left(\sup_{n}\frac{1}{n}\sum_{i=1}^{n}X_{i}^{p}\right)\int_{1}^{+\infty}x^{1-p}dx$, converges for $p>2$. Similarly, $\int_{1}^{+\infty}x\bar{F}^{\alpha}_{n}(x)dx$ converges for $p>\frac{2}{\alpha}.$
Therefore, $\hat{\xi}_{\alpha}^{w}(X)$ is a consistent estimator. This implies $\hat{\xi}_{\alpha}^{w}(X)\to\frac{(\alpha+4)}{6(\alpha+1)(\alpha+2)}$ under $H_{0}$. Hence the result. ∎
We can also develop uniformity test using $\hat{\xi}^{w}(X)$ which is given by
$\hat{\xi}^{w}(X)=-\frac{1}{2}\sum_{i=1}^{n-1}\left(X^{2}_{(i+1)}-X^{2}_{(i)}\right)\left(1-\frac{i}{n}\right)\log\left(1-\frac{i}{n}\right).$
Lemma 5.3.
Let $X_{1},X_{2},\cdots,X_{n}$ be a random sample drawn from a continuous distribution concentrated on [0,1]. Then, $0\leq\hat{\xi}^{w}(X)\leq\frac{1}{2e}$.
Proof.
The function $g(v)=-(1-v)\log(1-v),\;0<v<1,$ has maximum value $\frac{1}{2}$. Proceeding along the same line as in Lemma 5.1, we can get the desired result. ∎
Under $H_{0}$, $\xi^{w}(X)=\frac{5}{36}=0.1389$ which lies between $(0,\frac{1}{2e})$. Note that, $\frac{1}{2e}=0.1839$. So we can use $\hat{\xi}^{w}(X)$ as a test statistic for testing uniformity and we will reject the null hypothesis when either $\hat{\xi}^{w}(X)<C_{1}$ or $\hat{\xi}^{w}(X)>C_{2}$, where $C_{1}$ and $C_{2}$ will be calculated from the size condition. On using Theorem 7.1 of Mirali et al. (2017) we can see that, $\hat{\xi}^{w}(X)$ is consistent. So the uniformity test based on $\hat{\xi}^{w}(X)$ is also consistent.
Table 8: Power of the tests for various sample sizes.
$n$ Alternatives $\alpha=1$ $\alpha=2$ $\alpha=5$ $\alpha=7$ $\alpha=10$ ENT KS CvM AD
10 $A_{1.5}$ 0.1400 0.1742 0.1905 0.1683 0.1860 0.1420 0.1489 0.1660 0.1608
$A_{2}$ 0.3455 0.4228 0.4675 0.4609 0.4740 0.2893 0.3762 0.4378 0.4091
$B_{1.5}$ 0.0915 0.0865 0.0581 0.0456 0.0447 0.1884 0.0421 0.0330 0.0200
$B_{2}$ 0.2263 0.2055 0.1209 0.0907 0.0804 0.4363 0.0398 0.0226 0.0081
$B_{3}$ 0.5612 0.5549 0.3974 0.2917 0.2526 0.7958 0.0908 0.0486 0.0182
$C_{1.5}$ 0.1004 0.1191 0.1310 0.1222 0.1063 0.0302 0.1137 0.0978 0.1354
$C_{2}$ 0.1807 0.2184 0.2343 0.2063 0.1911 0.0330 0.2003 0.1491 0.2180
20 $A_{1.5}$ 0.2682 0.3230 0.3443 0.3468 0.3543 0.2411 0.2841 0.3184 0.3065
$A_{2}$ 0.6483 0.7555 0.8063 0.8105 0.8163 0.6207 0.7038 0.7726 0.7480
$B_{1.5}$ 0.2217 0.1739 0.0964 0.0762 0.0674 0.3039 0.0555 0.0502 0.0258
$B_{2}$ 0.5753 0.5012 0.2839 0.2014 0.1627 0.7115 0.1176 0.0972 0.1003
$B_{3}$ 0.9615 0.9466 0.8079 0.7058 0.5753 0.9905 0.4201 0.5089 0.5639
$C_{1.5}$ 0.1477 0.1732 0.1603 0.1516 0.1308 0.0589 0.1455 0.1224 0.1633
$C_{2}$ 0.3210 0.3897 0.3153 0.2806 0.2313 0.1461 0.3092 0.2538 0.3846
30 $A_{1.5}$ 0.3971 0.4644 0.5068 0.5357 0.5201 0.3268 0.3988 0.4713 0.4667
$A_{2}$ 0.8451 0.9204 0.9468 0.9506 0.9453 0.8201 0.8617 0.9192 0.9173
$B_{1.5}$ 0.3567 0.2768 0.1495 0.1202 0.0838 0.4047 0.0788 0.0585 0.0600
$B_{2}$ 0.8232 0.7506 0.4788 0.3880 0.2799 0.8849 0.2413 0.8044 0.3011
$B_{3}$ 0.9989 0.9968 0.9653 0.9263 0.8299 0.9991 0.7320 0.9969 0.9259
$C_{1.5}$ 0.2061 0.2567 0.2233 0.1803 0.1594 0.1112 0.1836 0.3184 0.2022
$C_{2}$ 0.4695 0.5798 0.4183 0.3313 0.2837 0.3577 0.4403 0.4016 0.5296
5.1 Critical points and Power of the tests
The exact distributions of $\hat{\xi}_{\alpha}^{w}(X)$ and $\hat{\xi}^{w}(X)$ are intractable. We use simulations to obtain the critical points for various choices of $\alpha$ and $n$. We generate 10000 random samples from U(0,1) distribution with different sample sizes and calculate the critical points at 5% level of significance. The critical points are provided in Table 7. In the table, by $\alpha=1$ we mean $\alpha\to 1$ i.e WCRE. We take $\alpha$ = 1, 2, 5, 7 and 10, and calculate the quantile points for $n$ = 10(5)50 and 50(10)100. For power calculation, we consider the following seven alternative distributions. The cdfs of these distributions are as follows:
$A_{j}:F(z)$ $=$ $1-(1-z)^{j},\;\;\;0\leq z\leq 1\;(j=1.5,\;2)$
[…]
$C_{j}:F(z)$ $=$ $\begin{cases}0.5-2^{j-1}(0.5-z)^{j},\;\;0\leq z\leq 0.5\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(j=1.5,\;2)\\ 0.5+2^{j-1}(z-0.5)^{j},\,\,0.5\leq z\leq 1\end{cases}$
These alternative distributions were first considered by Stephens (1974). The alternatives $A_{j}$’s provide points close to 0 than expected under uniformity and it is interpreted as a change in mean, the alternatives $B_{j}$’s provide points close to the middle i.e. 0.5 and it is interpreted as a change towards smaller variance, and finally, the alternatives $C_{j}$’s provide points towards both the extremes viz. 0 and 1, and it is interpreted as a shift towards larger variance. We compare power of our proposed tests with Kolmogorov-Smirnov (KS), Cramer-von Mises (CvM), Anderson-Darling (AD) tests and Vasicek sample entropy-based test of Dudewicz and Van Der Meulen (1981). The test statistic is defined as
$H_{mn}=\frac{1}{n}\sum_{k=1}^{n}\log[\frac{n}{2m}(X_{(k+m)}-X_{(k-m)})],$
where $m$ is the window size. Under $H_{0}$, $H_{mn}$ converges to zero and it is less than zero otherwise. So, we will reject $H_{0}$ when $H_{mn}$ is small. This test is called the ENT test. We compute the power for $n$ = 10, 20 and 30, and present the findings in Table 8. From Table 8
we observe that, ENT test of Dudewicz and Van Der Meulen (1981) performs better than other tests for alternatives $B_{j}$’s. Our proposed tests perform better than all the tests for alternatives $A_{j}$’s and $C_{j}$’s. When $\alpha$ = 2, proposed test performs better for alternative $C_{j}$’s and for higher values of $\alpha$ the proposed tests perform better for alternative $A_{j}$’s. Also note that, the WCRTE-based tests perfrom better than the test based on WCRE ($\alpha\to$1). The alternatives $A_{j}$’s and $C_{j}$’s provides points to the extreme. Our tests are based on weighted measures that not only consider the probabilistic information of the random variables but also the values that the random variables are taking. Thus, they are less influenced by the extreme observations. Therefore, we can rely on these measures when inference has to be made based on extreme observations and tail probabilities.
6 Conclusion
In this paper, we study some new properties of WCRTE measure and develop four non-parametric estimators of WCRTE. We investigate asymptotic properties of the proposed estimators as well. Also, we propose estimators for WCRE measure and study their asymptotic properties. We compare the performances of these estimators using simulation and develop two new uniformity tests. Power of these tests are compared with some popular existing tests and it is observed that, proposed tests perform very well compared to the other tests in majority of the cases.
We use one estimator of WCRTE and WCRE to construct the uniformity test. Other estimators may also be used for this purpose. It will be interesting to see for which value of the window size of these estimators give highest power. More work is needed in this direction.
Conflicts of Interest
The authors declare no conflict of interest.
Funding
No funding is received for this work.